Doing my homework can't finish working this problem; 4-3/5 + X = 5-1/10
did step 1) 23/5 + X = 51/10
step 2)???
don't know next steps, please help. Thank you in advance for your help. Candy

Answer by midwood_trail(310)

I assume you meant to type this:
4 - (3/5) + x = 5 - (1/10)
Multiply every term on both sides by the LCD, which is 10.
40 - 6 + 10x = 50 - 1
Can you take it from here?
If not, write back.